# Going beyond the creative limits of next-token prediction

### v3

@@ Appendix I Experiments on summarization
Experimental Details. In Table 3, we provide the hyperparameter details for the GPT models finetuned on both XSUM (Narayan et al., 2018) and CNN/DailyMail (Nallapati et al., 2016) for one epoch. We use a learning rate with linear warm up for $0.05$ of the total steps, followed by linear decay to 00. To measure Rouge and Self-Bleu, we generate and average across 5 summarizations per document, on a test dataset $T$ of $250$ datapoints. We finetune our models with either the NTP objective (Eq 2) or the teacherless MTP objective (Eq 3), with equal weight to both.
@@ I.2 Effect of seed-conditioning
One line of work relevant to us in the history of generative models is RNN-based VAE for text data (Bowman et al., 2016). The motivation, like in our work, was to learn high-level semantic features rather than next-token features with the hope of producing more novel sentences. However, this suffered from posterior collapse, where the model ignores the latent variable altogether inspiring various solutions (Yang et al., 2017; Goyal et al., 2017). Our results on noise-conditioning are also reminiscent of a line of work on exploration in reinforcement learning (RL), where it has been shown that adding noises to the policy model parameters enables more efficient exploration than directly adding noises to the output space (Plappert et al., 2017; Fortunato et al., 2017).
@@
Another theoretical question underlying generative models is that the optimum of their objectives are attained at perfect memorization; yet they tend to produce novel examples e.g., this question has been posed for GANs in Nagarajan et al. (2018) and for diffusion in Nakkiran et al. (2024) (see “remarks on generalization”) or Kamb & Ganguli (2024). Of relevance to us is, Kamb & Ganguli (2024) who provide a theoretical and empirical argument for how image diffusion models are able to generate combinatorially many creative outputs; theirs however do not require the type of planning our tasks do.

### v4

@@ Appendix I Experiments on summarization
Experimental Details. In Table 3, we provide the hyperparameter details for the GPT models finetuned on both XSUM (Narayan et al., 2018) and CNN/DailyMail (Nallapati et al., 2016) for one epoch. We use a learning rate with linear warm up for $0.05$ of the total steps, followed by linear decay to 0. To measure Rouge and Self-Bleu, we generate and average across 5 summarizations per document, on a test dataset $T$ of $250$ datapoints. We finetune our models with either the NTP objective (Eq 2) or the teacherless MTP objective (Eq 3), with equal weight to both.
@@ I.2 Effect of seed-conditioning
One line of work relevant to us in the history of generative models is RNN-based VAE for text data (Bowman et al., 2016). The motivation, like in our work, was to learn high-level semantic features rather than next-token features with the hope of producing more novel sentences. However, this suffered from posterior collapse, where the model ignores the latent variable altogether inspiring various solutions (Yang et al., 2017; Goyal et al., 2017). Our results on seed-conditioning are also reminiscent of a line of work on exploration in reinforcement learning (RL), where it has been shown that adding noise to the policy model parameters enables more efficient exploration than directly adding noises to the output space (Plappert et al., 2017; Fortunato et al., 2017).
@@
Another theoretical question underlying generative models is that the optimum of their objectives are attained at perfect memorization; yet they tend to produce novel examples e.g., this question has been posed for GANs in Nagarajan et al. (2018) and for diffusion in Nakkiran et al. (2024) (see “remarks on generalization”) or Kamb & Ganguli (2024). Okawa et al. (2023); Kamb & Ganguli (2024) study the creativity ability of image diffusion models in combining various concepts to create an image. Unlike our setting, these do not require planning.
Injecting input noise. We note that the concept of injecting input noise into a Transformer has been explored before, but for other functions (such as quality, robustness (Hua et al., 2022; Jain et al., 2024) or efficiency (Wang et al., 2024c).), and in other forms (e.g., inducing Gaussian noise).